Operation Summary for October 29, 2010 - November 1, 2010 |
The Integrated luminosity for the period from 10/25/10 to 11/1/10 was 59.99 pb-1. Friday Oct 29 The day shift began with Operations monitoring the Tevatron (TeV) store 8214 (213.98E30) and the Recycler stash (88.5E10), with the Antiproton Source (Pbar) stacking (13.2E10), with Linac, Booster, and Main Injector (MI) operating normally, with the NuMI experiment off for target work, with the MiniBooNE and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operators search and secured the NuMI enclosure and then NuMI personnel conducted low intensity target scans. Operations changed the antiproton stacking repetition rate to 2.4 seconds for the remainder of the weekend. The Neutron Therapy Facility (NTF) treated patients. During the evening shift, Operations terminated store 8214 at 4:07 PM. Operations prepared the CDF and D0 collision halls for access. NuMI completed their target scans at 5 PM. CDF and D0 came out of access at 7 PM. During the Tevatron turn on, a problem developed with the heater firing unit (HFU) at TeV sector D2; operators replaced the unit. Operations had the TeV on and ramping by 8:40 PM. The TeV orbits did not look good at the low beta magnets; TeV personnel came in to look into the problem. The investigation lasted until the end of shift. Operations disabled the T-978 beam at 6:06 PM. Saturday Oct 30 The midnight shift began with Operations monitoring the stash (385.7E10), with Pbar stacking (85.7E10), and with MiniBooNE taking beam. During the midnight shift, Operations put the TeV into HEP shot setup at 12:44 AM and established store 8219 at 1:57 AM with an average initial luminosity (AIL) of 312.69E30. Operations established beam to T-978 at 4:30 AM. Operators prepared NuMI for access at 6:45 AM. During the day shift, Operations terminated store 8219 at 11:30 AM, put the TeV into shot setup, and established store 8220 at 12:42 PM with an AIL of 341.6E30. Operators search and secured the NuMI target hall and then NuMI personnel began conducting low intensity target scans. Operations reported minor downtimes due to problems with the Linac, Booster, and MI RF systems. During the evening shift, NuMI completed their target scans at 5:23 PM. Operations disabled the T-978 beam at 6 PM. Sunday Oct 31 The midnight shift began with Operations monitoring store 8220 and the stash, with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8220 at 1:05 AM, put the TeV into shot setup, and established store 8221 at 2:10 AM with an AIL of 317.7E30. Operations established beam to T-978 at 4:20 AM. Operators prepared the NuMI target hall for access at 7 AM. Operations terminated store 8221 at 6:43 AM. Operators prepared the D0 collision hall for a 10-12 hour access at 7:22 AM. During the day shift, Operations monitored systems. During the evening shift, D0 came out of access at 5:48 PM. Operations turned on the TeV and checked the orbits. TeV personnel came in to help with orbit corrections. Operations prepared the CDF collision hall for access. Operations disabled the T-978 beam at 6 PM. Operations put the TeV into shot setup at 9:18 PM and established store 8227 at 10:12 PM with an AIL of 251.78E30. Operations reported minor downtimes due to problems with the Booster and Linac RF systems as well as the MI abort kickers. Monday Nov 1 The midnight shift began with Operations monitoring store 8227 and the stash, with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, Operations reported that the Pelletron started tripping off early on shift; Operations consulted with Electron Cooling (ECool) personnel. They concluded that the problem was inside the Pelletron and it would need to be accessed. Mechanical technicians came in and started to transfer the insulating gas (SF6) out of the Pelletron at 4 AM. Operations established beam to T-978 shortly after 4 AM. Operations reported minor downtimes due to Linac RF system trips. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator reported that Operations would start sending beam to NuMI after the next shot goes in. |
|
More Information Comments and Suggestions |
| last modified 11/01/2010 email Fermilab |
| Security, Privacy, Legal | |